OpenStudy (anonymous):

Find (f º g)(–2) if f(x) = 4x – 3 and g(x) = 3x + 3 Select one: a. 10 b. –4 c. –15 d. 0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (sidsiddhartha):

\[(f ºg)(x)=f[g(x)]\\use ~~this\]

OpenStudy (sidsiddhartha):

simply \[f[g(x)]=f[3x+3]=4(3x+3)-3=12x+12-3 \\=12x+9\]then \[f[g(-2)]-12*(-2)+9=-24+9=-15\] ok??

OpenStudy (anonymous):

so -15

OpenStudy (sidsiddhartha):

yeah ;)

OpenStudy (anonymous):

thanks can you help me on somemore please